# MUSE-Net: Missingness-aware mUlti-branching Self-attention Encoder for Irregular Longitudinal Electronic Health Records

### v1

@@ 3.2.3 Multi-branching Outputs
$\mathcal{L}(\boldsymbol{\omega};\hat{\mathcal{D}})=-\sum_{p=1}^{N}\sum_{i=1}^{N_{b}}I(p\in\mathcal{D}_{i})\left(l_{p}\log(\hat{P}^{(i)}(\hat{Y}_{p},\tilde{X}_{p};\boldsymbol{\omega}))+(1-l_{p})\log(1-\hat{P}^{(i)}(\hat{Y}_{p},\tilde{X}_{p};\boldsymbol{\omega}))\right)$ (12)
where $\boldsymbol{\omega}$ denotes the network parameters from both the self-attention encoder and the MB classifier; $I(\cdot)$ is an indicator function; $\hat{P}^{(i)}(\hat{Y}_{p},\tilde{X}_{p};\boldsymbol{\omega})$ is the predicted probability by output branch $i$ given the input data $\hat{Y}_{p}$ and corresponding missing value mask $\tilde{X}_{p}$. The final predicted probability for patient $p$ is computed as the average of the $N_{b}$ predicted probabilities: $\hat{P}_{MB}(\hat{Y}_{p},\tilde{X}_{p};\boldsymbol{\omega})=\sum_{i=1}^{N_{b}}\hat{P}^{(i)}(\hat{Y}_{p},\tilde{X}_{p};\boldsymbol{\omega})/N_{b}$, which is a more robust estimator compared to the single-branch (SB) counterpart as stated in Theorem I.
@@
$\bar{\hat{P}}^{(i)}(Y_{p})$ $=$ $\bar{\hat{P}}_{SB}(Y_{p}),~{}~{}~{}~{}~{}~{}~{}~{}~{}(i\in\{1,\dots,N_{b}\})$
$\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}^{(i)}(Y_{p})||\hat{P}^{(i)}(Y_{p};\boldsymbol{\omega}))]$ $=$ $\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{SB}(Y_{p})||\hat{P}_{SB}(Y_{p};\boldsymbol{\omega}))],$ (13)
@@
Theorem I shows that the MB classifier is more robust than the SB classifier (see Appendix 8.2 for the theoretical proof).
Figure 3: Illustration of the experimental design to evaluate the performance of the proposed MUSE-Net.
4 Experimental Design and Results
Fig. 3 shows our experimental design to evaluate the performance of the proposed MUSE-Net. We implement our MUSE-Net to investigate both a synthetic and real-world dataset for performance evaluation. The MUSE-Net is benchmarked with four popular model structures widely used in sequential modeling: LSTM [50], gated recurrent unit (GRU) [51], Time-aware LSTM (T-LSTM) [32], and TCN [30]. Additionally, we evaluate the prediction performance provided by different imputation methods, i.e., the proposed MGP+mask, MGP, single Gaussian process (GP), GP+mask, the baseline mean imputation, and Mean+mask. We also evaluate the impact of the number of branches on the final prediction performance. The dataset in both synthetic and real-world case studies is split into training (80%), validation (10%), and testing (10%) sets to facilitate a comprehensive training and evaluation process. The model performance is evaluated based on two key metrics: area under the receiver operating characteristic curve (AUROC) and area under the precision-recall curve (AUPRC). AUROC serves as a measure of the model’s ability to discriminate between classes, with a value of 1.0 indicating perfect classification and 0.5 denoting no discriminative power. AUPRC is especially valuable for assessing performance in classifying imbalanced datasets, reflecting the balance between precision and recall across various thresholds.
@@ 4 Experimental Design and Results
8:         if $label\_list[j]$ is 00 (negative instance) then $\triangleright$ For negative instance
9:             Generate $n\_variables$ time series data using $\text{ARMA}(p_{m},q_{m})$
10:         $\triangleright$ For positive instance
11:             Generate $n\_variables$ time series data using $\text{ARMA}(p_{m+n\_variables},q_{m+n\_variables})$
@@
4.1 Experimental Results in Synthetic Case Study
@@
Algorithm 1 shows the process to generate synthetic data from the autoregressive moving average (ARMA) model [52]. Specifically, we first generate three baseline time series variables by ARMA($p_{m}$,$q_{m}$) models with moving average of order $q_{m}$ and autoregressive of order $p_{m}$: $v_{m,t}=\phi_{m,1}v_{m,t-1}+\cdots+\phi_{m,p_{m}}v_{m,t-p_{m}}+\psi_{m,1}\epsilon_{m,t-1}+\cdots+\psi_{m,q_{m}}\epsilon_{m,t-q_{m}}+\epsilon_{m,t}$, where $\{\epsilon_{m,t}\}$ are uncorrelated random variables for variable $m$; $\Omega_{m}=\{\phi_{m,1},\cdots,\phi_{m,p_{m}},\psi_{m,1},\cdots,\psi_{m,q_{m}}\}$ is the set of coefficients for variable $m$. We initialize $\Omega_{m}$ randomly from a standard normal distribution and randomly select $p_{m}$,$q_{m}$ from set $\{1,2,3\}$. We generate the remaining seven variables through feature engineering applied to the three baseline time series, ensuring correlations exist among the variables and differences exist between different classes. Detailed information on the feature engineering process and the simulation parameters is available in the Appendix 8.1. The generated synthetic dataset is a 3-D tensor containing 5,000 samples with 90% from the majority class. Each sample has 10 variables measured across 50 time steps with irregular intervals. Notably, each variable is characterized by a distinct rate of missing values, varying randomly between 30% and 60%, mimicking real-world data complexities.
Fig. 4 shows the AUROC and AUPRC scores over training epochs for the MUSE-Net-9 (i.e., with 9 branching outputs) utilizing different imputation methods with missing value masks on the validation set of our simulated data. The ‘complete’ represents the scenario where MUSE-Net is trained on synthetic data without any missing values. The MUSE-Net-9+MGP method exhibits superior performance in both AUROC/AUPRC metrics and convergence speed. In contrast, the MUSE-Net-9+GP and MUSE-Net-9+mean methods display delayed convergence and lower scores, especially in the AUPRC metric.
Fig. 5 displays the AUROC and AUPRC scores across training epochs for MUSE-Net+MGP with different numbers of MB outputs on the simulated validation set. It is evident that within 8 epochs, all the MUSE-Nets achieve optimal performance, reaching near-perfect AUROC and AUPRC scores. This exceptional performance is partly due to the simplistic nature of the synthetic dataset and the high efficacy of MGP imputation and effective feature extraction achieved by the MUSE-Net. However, we observe variations in the convergence speed among different MB numbers. Specifically, as we increase the MB outputs from 1 to 9, the speed of convergence improves. This enhancement can be attributed to the improved balance in the training sub-dataset for each branching output. With only one output, the imbalanced ratio is as skewed as 9:1, presenting a significant challenge for the model to learn from the minority class. As the number of MB outputs increases to 9, the ratio between the positive and negative classes is approximately 1:1, providing a more balanced learning environment for each branch. As such, the network model can learn more effectively from both classes, leading to faster convergence.
@@
Figure 4: AUROC and AUPRC scores across epochs for MUSE-Net-9 utilizing different imputation methods on the validation set of the simulated data.
@@
Figure 5: AUROC and AUPRC scores across epochs for MUSE-Net with different numbers of MB outputs on the validation set of the simulated data imputed by MGP.
Figure 6: AUROC and AUPRC scores across epochs for MUSE-Net with mean imputation and different numbers of MB outputs on the validation set of the simulated data.
The impact of imbalanced data is more significant when modeling datasets with poor quality. Fig. 6 shows the AUROC and AUPRC scores across epochs for MUSE-Net with mean imputation and different numbers of MB outputs on the synthetic validation set. The mean imputation is known for introducing a high bias due to its simplistic assumption, resulting in a less reliable imputed dataset. According to Fig. 6, the MUSE-Net-9 converges within 6 epochs and achieves a good AUROC/AUPRC score, which indicates that our MB architecture together with the missing value masks effectively mitigates the negative impact of imbalanced datasets, even when the imputed training data is with low quality. In contrast, the MUSE-Net does not converge in 12 epochs if the MB architecture is not incorporated (i.e., MB 1) and the final AUROC/AUPRC scores decrease to 0.921/0.773 from 0.988/0.895.
4.2 Experimental Results in Real-world Case Study
We further conduct a real-world case study using the Diabetic Retinopathy (DR) dataset, obtained from the 2018 Cerner Health Facts data warehouse [7, 29], to evaluate our MUSE-Net. The medical variables selected for this study include 21 routine blood tests, 5 comorbidity indicators, 3 demographic variables, and the duration of diabetes. More details on variable selection can be referred to [29, 53]. Notably, the 21 blood tests are all subject to missing values, and the final dataset consists of records from 23,245 diabetic patients, with a minority of 8.9% diagnosed with DR. We use label “1” and “0” to indicate that a patient diagnosed with and without DR, respectively.
Fig. 7 shows the performance of our MUSE-Net compared to other benchmark models on the DR validation set. Note that to ensure a fair comparison, all models are designed to have similar sizes, with the number of model parameters approximately 4,300, as shown in Table 1. Furthermore, all models utilize the MGP imputation with missing value masks, 10 MB outputs, and Adam with decoupled weight decay optimizer [54]. According to Fig. 7, our MUSE-Net consistently outperforms other benchmarks, achieving the highest AUROC and AUPRC scores on the validation set. This trend is also evident in the results for the test set summarized in Table LABEL:tab:Model_scores_test, where MUSE-Net maintains the AUROC score of 0.949 and the AUPRC score of 0.883, which dominates those yielded by TCN, LSTM, T-LSTM, and GRU models. Specifically, the MUSE-Net has an improvement over TCN by 2.3% in AUROC and 5.3% in AUPRC, and an improvement over LSTM by 3.3% in AUROC and 7.1% in AUPRC, an improvement over T-LSTM by 2.4% in AUROC and 6.8%, and an improvement over GRU 1.9% in AUROC and 4.6% in AUPRC. This performance demonstrates that MUSE-Net is more effective at capturing the critical patterns inherent in irregular longitudinal DR data. The higher AUROC and AUPRC scores reflect the superior performance of our MUSE-Net in terms of both discrimination and precision-recall balance, which are critical for clinical decision making.
Table 1: Architectures of MUSE-Net and other benchmarks for DR dataset.
@@ 4.2 Experimental Results in Real-world Case Study
MUSE-Net 4,282 2 10 Adam with decoupled weight decay (AdamW) [54]
@@
Table 2: AUROC and AUPRC scores for MUSE-Net and other benchmarks on the DR test set.
MUSE-Net TCN LSTM T-LSTM GRU
AUROC 0.949 0.928 0.919 0.927 0.932
AUPRC 0.883 0.839 0.825 0.827 0.844
Table 3 presents the performance comparison in AUROC and AUPRC between our MUSE-Net-10 with MGP imputation and other imputation methods on the DR test set. It is worth noting that the MUSE-Net-10 with MGP+masks outperforms other methods, yielding an AUROC of 0.949 and an AUPRC of 0.883, which represents an improvement over GP+masks by 1.2% in AUROC and 6.1% in AUPRC, and an improvement over mean+masks by 3.5% in AUROC and 7.8% in AUPRC. Moreover, consistent performance improvement is achieved across all imputation methods when missing value masks are applied. Specifically, MGP+masks achieves an improvement over the non-mask MGP by 4.3% in AUROC and 8.5% in AUPRC. Similarly, employing masks with the GP imputation yields a 3.2% and 4.3% boost in AUROC and AUPRC respectively; the application of masks in mean imputation contributes to a 1.8% and 6.4% improvement in AUROC and AUPRC, respectively. The results show that the missing value masks enable the model to effectively account for the missingness patterns and mitigate potential discrepancies between imputed values and actual observations that may arise during the imputation process.
Table 3: The comparison of AUROC and AUPRC between our MUSE-Net-10 with MGP imputation and other imputation methods on the DR test set.
Model: MUSE-Net MGP imputation GP imputation Mean imputation
Mask w/o Mask Mask w/o Mask Mask w/o Mask
AUROC 0.949 0.910 0.938 0.909 0.917 0.901
AUPRC 0.883 0.814 0.833 0.799 0.819 0.770
Table 4 shows the AUROC and AUPRC scores for MUSE-Net with MGP imputation across various numbers of MB outputs in the DR test set. MB 1 corresponds to a model with a single output handling the original dataset with an imbalance ratio of approximately 1:10, and MB 10 represents a model with 10 outputs, each of which is trained with one of the 10 balanced sub-datasets with a ratio of approximately 1:1. This table suggests an overall increasing trend in both AUROC and AUPRC when the number of MB outputs increases. The highest AUROC/AUPRC scores of 0.949/0.883 are achieved with 10 MB outputs, demonstrating the enhanced predictive capability of our MUSE-Net in a balanced training environment. The performance gradually declines with a decrease in the number of MB outputs, reaching its lowest AUROC and AUPRC of 0.935 and 0.860 at MB 1. Specifically, the MUSE-Net with MB 10 achieves an improvement over the MUSE-Net with MB 1 by 1.5% in AUROC and 2.7% in AUPRC.
Table 4: AUROC and AUPRC scores for MUSE-Net+MGP with different MB numbers in the DR test set.
@@
AUROC 0.949 0.948 0.947 0.947 0.940 0.935
AUPRC 0.883 0.884 0.878 0.879 0.870 0.860
5 Conclusions

### v2

@@ III-B3 Multi-branching Outputs
$\mathcal{L}(\boldsymbol{\omega};\hat{\mathcal{D}})$ $=$ $-\sum_{p=1}^{N}\sum_{i=1}^{N_{b}}I(p\in\mathcal{D}_{i})\Big{(}l_{p}\log(\hat{P}^{(i)}(\hat{Y}_{p},\tilde{X}_{p};\boldsymbol{\omega}))$ (14)
$+(1-l_{p})\log(1-\hat{P}^{(i)}(\hat{Y}_{p},\tilde{X}_{p};\boldsymbol{\omega}))\Big{)}$
where $\boldsymbol{\omega}$ is the network parameters; $I(\cdot)$ is an indicator function; $\hat{P}^{(i)}(\hat{Y}_{p},\tilde{X}_{p};\boldsymbol{\omega})$ is the prediction by branch $i$ given the input $\hat{Y}_{p}$ and corresponding missing value mask $\tilde{X}_{p}$. The final predicted probability is computed as the average of the $N_{b}$ predictions: $\hat{P}_{MB}(\hat{Y}_{p},\tilde{X}_{p};\boldsymbol{\omega})=\sum_{i=1}^{N_{b}}\hat{P}^{(i)}(\hat{Y}_{p},\tilde{X}_{p};\boldsymbol{\omega})/N_{b}$, which is a more robust estimator compared to the single-branch (SB) counterpart as stated in Theorem I.
@@
$\bar{\hat{P}}^{(i)}(Y_{p})=\bar{\hat{P}}_{SB}(Y_{p}),~{}~{}~{}(i\in\{1,\dots,N_{b}\})$
$\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}^{(i)}(Y_{p})||\hat{P}^{(i)}(Y_{p};\boldsymbol{\omega}))]=$
$\mathbb{E}_{\boldsymbol{\omega}}[D_{KL}(\bar{\hat{P}}_{SB}(Y_{p})||\hat{P}_{SB}(Y_{p};\boldsymbol{\omega}))],$ (15)
@@
Theorem I shows that the MB classifier is more robust than the SB classifier (see Appendix VI-B for the theoretical proof).
IV Experimental Design and Results
We implement our MUSE-Net to investigate both a synthetic and real-world dataset for performance evaluation. The MUSE-Net is benchmarked with four models widely used in sequential data mining: LSTM [51], gated recurrent unit (GRU) [52], Time-aware LSTM (T-LSTM) [32], and TCN [53]. Additionally, we evaluate the prediction performance provided by different imputation methods, i.e., the proposed MGP+mask, MGP, single GP, GP+mask, the baseline mean imputation, and Mean+mask. We also evaluate the impact of the number of branches on the final prediction performance. The dataset is split into training (80%), validation (10%), and testing (10%) sets. The model performance is evaluated on four key metrics: area under the receiver operating characteristic curve (AUROC), area under the precision-recall curve (AUPRC), recall, and the F1 score. AUROC is a measure of the model’s ability to discriminate between classes. AUPRC is valuable for assessing performance in classifying imbalanced datasets, reflecting the balance between precision and recall. Recall assesses the model’s ability to correctly identify positive instances, while the F1 score provides a comprehensive evaluation by accounting for both the model’s sensitivity to positive cases and its overall predictive reliability.
@@ IV Experimental Design and Results
8:        if $label\_list[j]$ is 00 (negative instance) then
9:           Generate $n\_variables$ time series using $\text{ARMA}(p_{m},q_{m})$
10:        else
11:           Generate $n\_variables$ time series using $\text{ARMA}(p_{m+n\_variables},q_{m+n\_variables})$
@@
IV-A Experimental Results in Synthetic Case Study
@@
Algorithm 1 shows the process to generate synthetic data from the autoregressive moving average (ARMA) model [54]. Specifically, we first generate three baseline time series variables by ARMA($p_{m}$,$q_{m}$) models with moving average of order $q_{m}$ and autoregressive of order $p_{m}$: $v_{m,t}=\phi_{m,1}v_{m,t-1}+\cdots+\phi_{m,p_{m}}v_{m,t-p_{m}}+\psi_{m,1}\epsilon_{m,t-1}+\cdots+\psi_{m,q_{m}}\epsilon_{m,t-q_{m}}+\epsilon_{m,t}$, where $\{\epsilon_{m,t}\}$ are uncorrelated random variables for variable $m$; $\Omega_{m}=\{\phi_{m,1},\cdots,\phi_{m,p_{m}},\psi_{m,1},\cdots,\psi_{m,q_{m}}\}$ is the set of coefficients for variable $m$. We initialize $\Omega_{m}$ randomly from a standard normal distribution and randomly select $p_{m}$,$q_{m}$ from set $\{1,2,3\}$. We generate the remaining seven variables through feature engineering applied to the three baseline time series, ensuring correlations exist among the variables and differences exist between different classes. Detailed information on the feature engineering process and the simulation parameters is available in the Appendix VI-A. The generated synthetic dataset is a 3-D tensor containing 5,000 samples with 90% from the majority class. Each sample has 10 variables measured across 50 time steps with irregular intervals. Notably, each variable is characterized by a distinct rate of missing values, varying randomly between 30% and 60%, mimicking real-world data complexities.
@@
Figure 2: Evaluation scores across epochs for MUSE-Net-9 with different imputation methods on simulated validation data.
@@
Figure 3: Evaluation metrics scores across epochs for MUSE-Net with varying MB outputs on MGP-imputed validation data.
Fig. 2 shows the AUROC, AUPRC, F1 score, and Recall on validation dataset over training epochs for MUSE-Net-9 (i.e., with 9 branching outputs) using different imputation methods with missing value masks on the validation set of our simulated data. The ’complete’ scenario represents the case where MUSE-Net is trained on synthetic data without any missing values. Among all methods, MUSE-Net-9+MGP demonstrates the best performance across all four metrics, achieving higher AUROC, AUPRC, F1 score, and Recall, along with the fastest convergence. In contrast, the MUSE-Net-9+GP and MUSE-Net-9+mean methods exhibit slower convergence and lower overall performance, highlighting their limitations in handling missing values effectively.
Fig. 3 presents the AUROC, AUPRC, F1 score, and Recall across training epochs for MUSE-Net+MGP with different numbers of MB outputs on the simulated validation set. Models with a higher number of MB outputs (MB5, MB7, and MB9) demonstrate faster convergence and superior overall performance. This improvement can be attributed to the enhanced class balance within each branching output. When only one MB output is used (MB1), the class imbalance is severe (approximately 9:1), making it difficult for the model to effectively learn from the minority class. In contrast, increasing the number of MB outputs to 9 results in a more balanced class ratio (1:1) for each branch, facilitating better learning from both classes.
The F1 score and Recall curves further emphasize the differences in learning dynamics. Models with more MB outputs exhibit rapid improvements in these metrics, while those with fewer MB outputs (MB1 and MB3) show delayed growth and lower overall performance, particularly in the early training stages. This suggests that models with fewer MB outputs struggle to effectively capture minority class patterns, leading to slower convergence and reduced recall. Overall, increasing the number of MB outputs enhances both convergence speed and classification performance, reinforcing the benefits of a more balanced learning framework.
TABLE I: Architectures of MUSE-Net and other benchmarks for DR dataset.
@@
MUSE-Net 4,010 2 10 Adam with decoupled weight decay (AdamW) [55]
@@
IV-B Experimental Results in Real-world Case Study
We further conduct a case study using the Diabetic Retinopathy (DR) dataset, obtained from the 2018 Cerner Health Facts data warehouse [7, 30], to evaluate our MUSE-Net. The variables selected for this study include 21 routine blood tests, 5 comorbidity indicators, 3 demographic variables, and the duration of diabetes. More details on variable selection can be referred to [30, 56]. Notably, the 21 blood tests are all subject to missing values, and the final dataset consists of records from 23,245 diabetic patients, with a minority of 8.9% diagnosed with DR. We use label “1” and “0” to indicate that a patient diagnosed with and without DR, respectively.
Fig. 4 shows the performance of MUSE-Net compared to other benchmark models on the DR validation set. Note that to ensure a fair comparison, all models are designed to have similar sizes, with the number of model parameters approximately 4,000, as shown in Table I. Furthermore, all models utilize the MGP imputation with missing value masks, 10 MB outputs, and Adam with decoupled weight decay optimizer [55]. According to Fig. 4, our MUSE-Net consistently outperforms other benchmarks, achieving the highest AUROC and AUPRC on the validation set. This trend is also evident in the results for the test set summarized in Table LABEL:tab:Model_scores_test, where MUSE-Net maintains the AUROC of 0.949 and the AUPRC of 0.883, which dominates those yielded by TCN, LSTM, T-LSTM, and GRU models. Specifically, the MUSE-Net has an improvement over TCN by 2.3% in AUROC and 5.3% in AUPRC, and an improvement over LSTM by 3.3% in AUROC and 7.1% in AUPRC, an improvement over T-LSTM by 2.4% in AUROC and 6.8%, and an improvement over GRU 1.9% in AUROC and 4.6% in AUPRC. This performance demonstrates that MUSE-Net is more effective at capturing the critical patterns inherent in irregular longitudinal DR data.
TABLE II: AUROC, AUPRC, Recall and F1 scores for MUSE-Net and other benchmarks on the DR test set.
MUSE-Net TCN LSTM T-LSTM GRU
AUROC 0.952 0.935 0.919 0.932 0.932
AUPRC 0.886 0.848 0.825 0.844 0.844
F1 0.750 0.704 0.580 0.671 0.671
Recall 0.850 0.809 0.836 0.827 0.827
@@ IV-B Experimental Results in Real-world Case Study
Table III presents the performance comparison in AUROC and AUPRC between our MUSE-Net-10 with MGP imputation and other imputation methods on the DR test set. It is worth noting that the MUSE-Net-10 with MGP+masks outperforms other methods, yielding an AUROC of 0.952 and an AUPRC of 0.886, which represents an improvement over GP+masks by 2.0% in AUROC and 7.1% in AUPRC, and an improvement over mean+masks by 2.5% in AUROC and 1.5% in AUPRC. Moreover, consistent performance improvement is achieved across all imputation methods when missing value masks are applied. Specifically, MGP+masks achieves an improvement over the non-mask MGP by 5.7% in AUROC and 10.6% in AUPRC. The results show that the missing value masks enable the model to effectively account for the missingness patterns and mitigate potential discrepancies between imputed and actual values that may arise during the imputation process.
TABLE III: AUROC and AUPRC comparison of MUSE-Net-10 with MGP and other imputation methods on the DR test set.
Model: MUSE-Net-10 MGP GP Mean
Mask w/o Mask w/o Mask w/o
AUROC 0.952 0.901 0.933 0.889 0.929 0.885
AUPRC 0.886 0.801 0.827 0.791 0.876 0.789
Table IV shows the AUROC and AUPRC scores for MUSE-Net with MGP imputation across various numbers of MB outputs in the DR test set. MB 1 corresponds to a model with a single output handling the original dataset with an imbalance ratio of approximately 1:10, and MB 10 represents a model with 10 outputs, each of which is trained with one of the 10 balanced sub-datasets with a ratio of approximately 1:1. This table suggests an overall increasing trend in both AUROC and AUPRC when the number of MB outputs increases. The highest AUROC/AUPRC scores of 0.952/0.886 are achieved with 10 MB outputs, demonstrating the enhanced predictive capability of our MUSE-Net in a balanced training environment. The performance gradually declines with a decrease in the number of MB outputs, reaching its lowest AUROC and AUPRC of 0.940 and 0.864 at MB 1.
TABLE IV: AUROC and AUPRC scores for MUSE-Net+MGP with different MB numbers in the DR test set.
@@
AUROC 0.952 0.947 0.946 0.947 0.945 0.940
AUPRC 0.886 0.876 0.877 0.886 0.869 0.864
IV-C Intepretability Analysis of MUSE-Net in DR Prediction
We further analyze the attention weights learned by the interpretable multi-head attention mechanism to interpret the model’s decision-making process for DR prediction.
IV-C1 Averaged Attention Map Across Layers
Figure 5: The averaged attention maps over all test samples for the first and second layers of the MUSE-Net: (a) the first layer; (b) the second layer
Fig. 5 presents the averaged attention maps over all test samples for the first and second layers of the MUSE-Net. Each heatmap represents the attention weight distribution between query (y-axis) and key time steps (x-axis), where higher values indicate stronger attention. These maps help clinicians interpret how the model makes predictions by identifying which time steps play a dominant role in decision-making. In the first layer (Fig. 5(a)), the attention distribution is skewed toward the later time, with the highest attention at the final time step. This suggests that the model places greater importance on more recent time steps when making predictions. Earlier time steps receive lower attention, indicating that they contribute less to the decision-making in the initial processing layer. Clinically, this aligns with the understanding that recent variable changes carry the most relevant information for assessing DR progression.
In the second layer (Fig. 5(b)), attention weights are more evenly distributed. This indicates that, after the first layer refines feature representations, the second layer integrates a more comprehensive view of temporal dependencies, balancing both recent and earlier observations. From a clinical perspective, this suggests that while initial assessments focus on recent trends, a deeper analysis incorporates long-term history to ensure a comprehensive decision-making process. This hierarchical mechanism mirrors how clinicians evaluate patient histories: first prioritizing the most recent indicators, then considering long-term trends for a more thorough assessment.
IV-C2 Attention Patterns in DR and Non-DR
Fig. 6 shows the column-wise sum of attention weights across all time steps for DR and Non-DR patients, providing insight into how much total attention is allocated to each time step for different groups in the test dataset. Fig. 6(a) shows the sum of attention weights in the first layer, revealing a clear difference: DR samples exhibit a more evenly distributed attention pattern across time steps. This suggests that the model considers multiple time steps when identifying DR, which is clinically relevant because DR progression is a gradual process influenced by long-term trends of many risk factors. Non-DR samples, on the other hand, show a steady increase in attention allocation over time, meaning that recent observations are more predictive of non-DR. This aligns with clinical practice where the absence of warning signs in recent medical history reduces the disease likelihood. This divergence in attention distribution provides critical insights: for DR patients, historical data remains important in assessing disease progression, whereas for non-DR patients, absence of recent warning signs is a stronger predictor of continued healthy status.
Fig. 6(b) presents the summed attention weights from the second layer. Here, both DR and non-DR samples follow an increasing trend, but a key observation is that the scale of the y-axis is smaller than in the first layer, indicating that attention is more evenly distributed. This suggests that each time step retains its importance, ensuring that no single time step dominates the decision-making process for both DR and Non-DR patients. These findings demonstrate that MUSE-Net not only achieves strong predictive performance but also offers clinically interpretable insights by mimicking how human experts evaluate temporal patient data in DR diagnosis.
Figure 6: The column-wise sum of attention weights across all time steps in the test dataset: (a) the first layer; (b) the second layer
V Conclusions